OpenStudy (anonymous):

Can you explain how to find the x-intercept in the equation: y=x^2-12x+39 In detail pleaaasse/

OpenStudy (anonymous):

The graph of this relation crosses the x-axis when y = 0. That is, when x^2 - 12x + 39 = 0. making sense so far?

OpenStudy (jamesj):

I think there's a problem there.

hero (hero):

Yeah, it's complex

OpenStudy (jamesj):

In your calculations: b^2 - 4ac = 144 - 156 < 0

hero (hero):

As you can see I figured it out already

hero (hero):

I didn't do it the conventional way though. I'm just saying that because you assumed that I used the quad formula

OpenStudy (jamesj):

lol. After I pointed out to you there was a problem. But ok.

hero (hero):

It's all good James.

OpenStudy (anonymous):

Who cares what method you use. If b^2 - 4ac is negative, there are no real roots.

hero (hero):

Simon, why do even comment on the matter? I didn't use the quad formula at all.

hero (hero):

No part of it